OpenStudy (anonymous):

If the hypotenuse of a right triangle is 2 sqrt 3 units long and one of the legs is sqrt 3 units long, then how long is the other leg? 3 2 sqrt 3 sqrt 3

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (here_to_help15):

And you dont call me lol

OpenStudy (anonymous):

sqrt 3?

OpenStudy (here_to_help15):

f(2sqrt(3))^2 = (sqrt(3))^2 + x^2 12 = 3 + x^2 9 = x^2 x + 3

OpenStudy (anonymous):

so its 3?

OpenStudy (here_to_help15):

Yea/

OpenStudy (anonymous):

thnx